OpenStudy (anonymous):

use elimination to slove this system of equations. 3y=2x+1 -6y=-6x-12 A.(7,5) B.(-5,7) C.(5,7) D.(7,6)

OpenStudy (anonymous):

y = 2x + 1 / 3 plug in second equation

OpenStudy (turingtest):

that would be substitution, not elimination

OpenStudy (anonymous):

ohh oops sorry dont rememebr the methods name

OpenStudy (turingtest):

yeah, the names are weak in my opinion... to use elimination, multiply the first equation by 2 and add it to the second. that will "eliminate" y

OpenStudy (anonymous):

i got a weird answer.

OpenStudy (turingtest):

can you show me what you got after just the first step I said?

OpenStudy (anonymous):

6y=4x+22 -6y=-6x-12

OpenStudy (turingtest):

ok, now add those two equations together what do you get?

OpenStudy (anonymous):

-2x+10?

OpenStudy (turingtest):

oh you have an extra 2 in there, should be 6y=4x+2 -6y=-6x-12 unless there was a typo in your original post which is it?

OpenStudy (anonymous):

extra 2

OpenStudy (turingtest):

6y=4x+2 or 6y=4x+22 ???

OpenStudy (turingtest):

what are your two original equations again?

OpenStudy (anonymous):

i got 6y=4x+22 3y=2x+1 -6y=-6x-12 are the original

OpenStudy (turingtest):

ok, so multiplying the first equation by 2 how do you get 22 ?

OpenStudy (turingtest):

multiply the first equation by 2 and you get 2(3y=2x+1) ^^^^ simplify, what do you get?

OpenStudy (anonymous):

-2x+3y=1 6x-6y=-12 6(-2x+3y=1) -2(6x-6y=-12) =-12x+18y=6 ----- i and -12x+12y=-24 -----ii subtractin i and ii we get 6y=30 therefore y=30/6=5 putting y=5 in -2x+3y=1 we get -2x+15=1 x=14/-2 therefore x=-7 ans. (-7,5)